# Coupled cluster method tailored with quantum computing

### v1

@@ I Introduction
Active spaces are designed to consider only the most important orbitals for treating the strong correlation of electrons (i.e., static correlation). The active spaces, however, naturally ignore the remaining weak correlation (i.e., dynamical correlation) resulting from the non-active orbitals. The active space approximation has been extensively discussed and addressed in traditional quantum chemistry [4, 5, 6, 7, 8, 9, 10].
In quantum computing, various approaches have been proposed to incorporate ignored electron correlation [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. These approaches fall into two broad categories. The first involves constructing models incorporating weak electron correlation during the development of effective Hamiltonians, which are then solved using quantum computers. This method is often called the ‘perturb-then-diagonalize’ approach. The second category improves the results based on an active space Hamiltonian solution obtained by quantum computers and is known as the ‘diagonalize-then-perturb’ approach. The former requires an up-front estimation of the electron correlation outside the active space, which introduces some arbitrariness. The latter commonly uses the internally-contracted multireference perturbation theory, or multireference configuration interaction, and there have already been proposals for implementation using quantum computing results. However, these classical computing methods require higher-order reduced density matrices (RDMs), leading to prohibitive measurement costs on quantum computers and seems impractical [34, 35, 22]. Consequently, there is an increasing demand for methods considering weak electron correlation without relying on higher-order RDMs.
@@
This study proposes combining TCC and quantum computing to recover dynamical correlation from excluded orbitals outside the active space. Our practical method is a two-step approach using a quantum computer to solve the static correlation and add the remaining dynamical correlation using classical computers. This method requires only a limited number of qubits to mainly solve the static correlation, while the relatively computationally cheap consideration of weak correlation can be performed on classical hardware. To extract the solved quantum state from a quantum computer, we use a tomographic method, computational basis tomography (CBT) [57]. CBT is based on computational basis sampling [57] and is an effective strategy when only a small number of Slater determinants are non-negligible, as is often the case in quantum chemistry. Additionally a method is proposed, in which we utilize the direct access to the energy of the active space and add our newly determined electron correlation outside the active space. The accuracy of the proposed method is further enhanced using this approach.
@@ II.1 Tailored Coupled Cluster
where $\hat{T}_{i}^{a}$ and $\hat{T}_{i,j}^{a,b}$ are the single and double excitation CC operators. Say, $\hat{T}_{i,j}^{a,b}$ excites electrons from the occupied orbitals $i$ and $j$ to the unoccupied orbitals $a,b$. The brackets {} in Eq. (3) indicates that at least one of the $i,j,a,b$ orbitals is not present in the active space.
In the first step of TCC, the molecule is described using a variational method using an active space approximation, such as the CASCI approach. This step aims to incorporate strong correlation by assuming that all the strongly correlated degrees of freedom are involved in the active space. Electron configurations’ coefficients of the varitational wavefunction |$\ket{\psi^{\text{active}}_{\text{{variational}}}}$ can be mapped directly to coupled cluster amplitudes through the known relationship between their corresponding coupled cluster and configuration interaction (CI) operators:
@@
where $\hat{C}_{1}$ and $\hat{C}_{2}$ are the CI operators to create single and double excitations, respectively. This procedure allows us to approximately reconstruct the variational wavefnction |$\ket{\psi^{\text{active}}_{\text{{variational}}}}$ in the CCSD ansatz:
@@
Additionally, given that TCC is a method rooted in CCSD, after optimizing the CCSD operator $e^{\hat{T}^{\text{{rest}}}(\theta^{\text{{rest}}})}$, one can further improve accuracy by perturbatively incorporating the effects of the three-body excitation operator in the same way as in CCSD(T). So we denote TCC with the (T) correction as TCC(T) in this study. For TCC(T), all active single and double amplitudes must be set to zero during the (T) correction calculation to prevent double-counting [38, 58].
@@ II.2 Computational Basis Tomography
where $1/\mathcal{N}$ is the normalization constant and $|k_{i}\rangle$ is a computational basis (i.e., Slater determinants or electron configuration) listed in the order of the absolute values of the coefficients $\langle k_{i}|\Psi\rangle$. The coefficients $\langle k_{i}|\Psi\rangle$ are equivalent to the CI coefficients. The parameter $R$ is introduced to truncate the trivial computational basis; therefore, $R$ needs to be chosen large enough in order to adequately approximate the original quantum state $|\Psi\rangle$.
@@
To determine the phase factor we need to measure the interference factors $\langle k_{i}|\Psi\rangle\langle\Psi|k_{j}\rangle$. In the case $k_{i}=k_{j}$ this is equivalent to a determine $|\langle k_{i}|\Psi\rangle|^{2}$. When $k_{i}\neq k_{j}$ we can rewrite the interference factor as
@@
$\langle k_{i}|\Psi\rangle\langle\Psi|k_{j}\rangle=\frac{(\langle k_{1}|\Psi\rangle\langle\Psi|k_{i}\rangle)^{*}\langle k_{1}|\Psi\rangle\langle\Psi|k_{j}\rangle}{|\langle k_{1}|\Psi\rangle|^{2}}.$ (12)
@@ II.3 QC-CBT-TCC
In the first step we determine the active-space ground state |$\ket{\psi_{\text{QC}}^{\text{active}}}$ using a quantum computer. Different quantum (or quantum-classical hybrid) algorithms, such as variational quantum eigensolver (VQE) [60, 61], quantum phase estimation (QPE) [62], and qunatum imaginary time evolution [63, 64, 65, 66, 67] can be used in this step.
@@
The QC-CBT-TCC energy $E_{\textrm{QC-CBT-TCC}}$ is then given by
@@ II.4 Enhanced QC-CBT-TCC
One of the issues associated with QC-CBT-TCC is the lack of assurance that the energy obtained via QC-CBT-TCC will outperform classical algorithms in accuracy. This arises because of the approximate nature of both CBT and TCC methodologies. To address this problem, we introduce a correction method as described below. The corrected energy $E_{\rm QC-CBT-TCC(c)}$ is defined as:
@@
is the CCSD energy after dressing the CCSD amplitudes but before the remaining CCSD optimization. In that case, the remaining operators $e^{\hat{T}^{\text{rest}}}$ are equivalent to the identity.
Note that Izsák et al. has already employed a similar extrapolative correction method in Ref. [68] in the context of using quantum computers for quantum chemistry as well as Daniel Kats and his co-workers for FCIQMC-TCC [43].
@@ III Results and Discussion
The computational details are as follows. A chemqulacs [69] library was used to simulate quantum circuits on classical hardware. We employed disentangled unitary coupled cluster singles and doubles (UCCSD) as an ansatz for VQE and BFGS as an optimizer. CCSD, CCSD(T), CASCI, and FCI were calculated using PySCF [70]. The MR-CISD+Q method was employed to obtain reference data for the PECs of H2O and N2 because the FCI calculation was too computationally demanding for the hardware at hand. This method was implemented in the ORCA quantum chemistry program Version 5.0.3 [71]. All calculations for the PECs in this study were performed using the cc-pVDZ basis set. In Section III.3, we used the 6-31G*** basis set to compute the activation energy for the Cope rearrangement. We used the canonical orbitals for all molecules except for LiH, which was calculated using the natural orbitals provided by a CCSD calculation. This, however, is an approximation because the CCSD implementation of PySCF used in this study assumes canonical orbitals, so off-diagonal terms in the Fock matrix are ignored. The active spaces were determined using the lowest energy level unoccupied orbitals and the highest energy occupied orbitals from the Hartree-Fock calculation. To transfer the CI coefficients determined by CBT to the amplitudes for CCSD only the real parts were considered. This was necessary because usually the CCSD for a non-relativistic or non-periodic Hamiltonian expects real coefficients. The CI coefficients were normalized to ensure the normalization. Unless otherwise indicated, we applied the parameters $R=100$, $N_{\text{sample}}=10^{6}$, $N_{U}=10^{6}$, and $N_{V}=10^{6}$ for all the CBT calculations.
@@ III.1 Potential energy curves
First, we investigated the PEC of the LiH molecule. In addition to the QC-CBT-TCC(c), we show the results for QC-CBT-TCC, active space UCCSD, and FCI, as well as HF, CCSD, and CCSD(T). The selected active spaces for these calculations were two electrons and two spacial orbitals, the highest occupied and the lowest unoccupied orbitals.
Figure 2 (a) shows the PEC of the LiH molecule. All considered methods followed the qualitative features of the FCI energy curve shape, except for HF. The active space UCCSD increased the accuracy compared to that of the HF solution. This behavior is most notable in the bond dissociation region. At the equilibrium bond length, the active space UCCSD can capture some dynamical correlation. The QC-CBT-TCC and its enhanced version, the QC-CBT-TCC(c), further increased the computed energy’s accuracy. This indicates that additional dynamical correlations outside the active space are required to reach good ground-state energy. CCSD and CCSD(T) both gave energies very similar to FCI. In the high dissociation regime, where static correlation dominates, all methods except for HF gave similar results.
We focus on the energy errors of the coupled cluster and present methods compared with the FCI energy in Figure 2 (b). CCSD and CCSD(T) showed better accuracy throughout the tested bond lengths than the QC-CBT-TCC and QC-CBT-TCC(c). This can be attributed to the small static correlation within the molecule. For such a system, CCSD and CCSD(T) can capture electron correlation up to high accuracy. The energy difference between the QC-CBT-TCC and QC-CBT-TCC(c) was small. However, the QC-CBT-TCC(c) appears to produce a smoother energy error. We assume that some errors in the QC-CBT-TCC did cancel out when the QC-CBT-TCC(c) is constructed.
@@
In Figure 3 (b), we show the explicit error of the presented methods compared to MR-CISD+Q. CCSD(T), and arguably CCSD also break in the highly entangled region. QC-CBT-TCC and the QC-CBT-TCC(c) can give accurate energies. However, in the high dissociation region, QC-CBT-TCC’s error increases, whereas that of QC-CBT-TCC(c) shows this effect is reduced, indicating that the QC-CBT-TCC(c) method produces more stable energies.
@@ III.2 Number of shot dependency of QC-CBT-TCC
CBT used in our method adds a certain level of uncertainty to our computed energies. Determining the CASCI coefficients using CBT is a statistical process. To investigate this statistical behavior, we examined for the molecules LiH, H2O, and N2 the impact of varying measurement repetitions in CBT on the QC-CBT-TCC(c). We considered two constellations, one for the molecules at the equilibrium bond length [36] and one for a bond length of twice the equilibrium bond distance. Note that the geometries were not optimized. To ensure statistical significance, we repeated each measurement setup 1000 times and showed its effect on the QC-CBT-TCC(c). The parameter $R$, used to truncate the trivial computational basis in CBT, is set as 100. The behavior of LiH was similar to that of H2O and N2, and the results shown in the Appendix. We demonstrated the effect using a box blot to demonstrate numerical data graphically. Whiskers represent the spread of the recorded data, excluding outliers represented as single dots. The box indicates the range from the first to the third quartiles, with the median drawn in the middle.
@@
Figure 5: Energy insecurity of QC-CBT-TCC(c) for H2O with a different number of measurements for CBT. $N_{\text{sample}},N_{U}$, and $N_{V}$ were all measured in the specified quantities in the figure as the X-axis values. CBT was performed 1000 times for each setting to obtain an estimate for the statistical errors. (a) Number of measurements dependency at an equilibrium distance of 1.808 Bohr. (b) Number of measurements dependency at 3.617 Bohr, twice the equilibrium distance.
@@
Figure 6: Energy insecurity of QC-CBT-TCC(c) for N2 with a different number of measurements for CBT. $N_{\text{sample}},N_{U}$, and $N_{V}$ were all measured in the specified quantities in the figure as the X-axis values. CBT was performed 1000 times for each setting to obtain an estimate for the statistical errors. (a) Number of measurements dependency at an equilibrium distance of 2.060 Bohr. (b) Number of measurements dependency at 4.119 Bohr, twice the equilibrium distance.

### v2

@@ I Introduction
Active spaces are designed to consider only the most important orbitals for treating the strong correlation of electrons (i.e., static correlation). The active spaces, however, naturally ignore the remaining weak correlation (i.e., dynamical correlation) resulting from the non-active orbitals. Moreover, as the selection of an active space is more or less based on the chemical intuition of the user, there is often a degree of arbitrariness. This user dependency can have a significant impact on the accuracy of the methods using an active space. Recently automated active space selection procedures have been proposed to reduce user bias and improve the reliability of results [4, 5, 6, 7, 8, 9, 10, 11]. The active space approximation has been extensively discussed and addressed in traditional quantum chemistry [12, 13, 14, 15, 16, 17, 18].
In quantum computing, various approaches have been proposed to incorporate ignored electron correlation [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. These approaches fall into two broad categories. The first involves constructing models incorporating weak electron correlation during the development of effective Hamiltonians, which are then solved using quantum computers. This method is often called the ‘perturb-then-diagonalize’ approach. The second category improves the results based on an active space Hamiltonian solution obtained by quantum computers and is known as the ‘diagonalize-then-perturb’ approach. The former requires an up-front estimation of the electron correlation outside the active space, which introduces some arbitrariness. The latter commonly uses the internally-contracted multireference perturbation theory or multireference configuration interaction, and there have already been proposals for implementation using quantum computing results. However, these classical computing methods require higher-order reduced density matrices (RDMs), leading to prohibitive measurement costs on quantum computers and seem impractical [42, 43, 30, 44]. Consequently, there is an increasing demand for methods considering weak electron correlation without relying on higher-order RDMs.
@@
This study proposes combining TCC and quantum computing to recover dynamical correlation from excluded orbitals outside the active space. Our practical method is a two-step approach using a quantum computer to solve the static correlation and add the remaining dynamical correlation using classical computers. This method requires only a limited number of qubits to mainly solve the static correlation, while the relatively computationally cheap consideration of weak correlation can be performed on classical hardware. To extract the solved quantum state from a quantum computer, we use a tomographic method, computational basis tomography (CBT) [66]. CBT is based on computational basis sampling [66] and is an effective strategy when only a small number of Slater determinants are non-negligible, as is often the case in quantum chemistry. Additionally, a method is proposed, in which we utilize the direct access to the energy of the active space and add our newly determined electron correlation outside the active space. The accuracy of the proposed method is further enhanced using this approach.
@@ II.1 Tailored Coupled Cluster
where $\hat{T}_{i}^{a}$ and $\hat{T}_{i,j}^{a,b}$ are the single and double excitation CC operators. Say, $\hat{T}_{i,j}^{a,b}$ excites electrons from the occupied orbitals $i$ and $j$ to the unoccupied orbitals $a,b$. The brackets {} in Eq. (3) indicate that at least one of the $i,j,a,b$ orbitals is not present in the active space.
In the first step of TCC, the molecule is described using a variational method using an active space approximation, such as the CASCI approach. This step aims to incorporate strong correlation by assuming that all the strongly correlated degrees of freedom are involved in the active space. Electron configurations’ coefficients of the variational wavefunction |$\ket{\psi^{\text{active}}_{\text{{variational}}}}$ can be mapped directly to coupled cluster amplitudes through the known relationship between their corresponding coupled cluster and configuration interaction (CI) operators:
@@
where $\hat{C}_{1}$ and $\hat{C}_{2}$ are the CI operators to create single and double excitations, respectively. This procedure allows us to approximately reconstruct the variational wavefunction|$\ket{\psi^{\text{active}}_{\text{{variational}}}}$ in the CCSD ansatz.
@@
Additionally, given that TCC is a method rooted in CCSD, after optimizing the CCSD operator $e^{\hat{T}^{\text{{rest}}}(\theta^{\text{{rest}}})}$, one can further improve accuracy by perturbatively incorporating the effects of the three-body excitation operator in the same way as in CCSD(T). So, we denote TCC with the (T) correction as TCC(T) in this study. For TCC(T), all active single and double amplitudes must be set to zero during the (T) correction calculation to prevent double-counting [47, 67].
@@ II.2 Computational Basis Tomography
where $1/\mathcal{N}$ is the normalization constant and $|k_{i}\rangle$ is a computational basis (i.e., Slater determinants or electron configuration) listed in the order of the absolute values of the coefficients $\langle k_{i}|\Psi\rangle$. The coefficients $\langle k_{i}|\Psi\rangle$ are equivalent to the CI coefficients. The parameter $R$ is introduced to truncate the trivial computational basis; therefore, $R$ needs to be chosen large enough in order to adequately approximate the original quantum state $|\Psi\rangle$. We tentatively consider the upper limit of $R$ to be around 1000-10000. By keeping the number of extracted coefficients $R$ constant, the method sacrifices some accuracy but enables scalability to large systems.
@@
To determine the phase factor, we need to measure the interference factors $\langle k_{i}|\Psi\rangle\langle\Psi|k_{j}\rangle$. In the case $k_{i}=k_{j}$ this is equivalent to a determine $|\langle k_{i}|\Psi\rangle|^{2}$. When $k_{i}\neq k_{j}$, we can rewrite the interference factor as
@@
$\langle k_{i}|\Psi\rangle\langle\Psi|k_{j}\rangle=\frac{(\langle k_{1}|\Psi\rangle\langle\Psi|k_{i}\rangle)^{*}\langle k_{1}|\Psi\rangle\langle\Psi|k_{j}\rangle}{|\langle k_{1}|\Psi\rangle|^{2}},$ (12)
@@ II.3 QC-CBT-TCC
In the first step, we determine the active-space ground state |$\ket{\psi_{\text{QC}}^{\text{active}}}$ using a quantum computer. Different quantum (or quantum-classical hybrid) algorithms, such as variational quantum eigensolver (VQE) [69, 70], quantum phase estimation (QPE) [71], and qunatum imaginary time evolution [72, 73, 74, 75, 76] can be used in this step.
@@
The QC-CBT-TCC energy $E_{\textrm{QC-CBT-TCC}}$ is then given by the following projected energy:
@@ II.4 Enhanced QC-CBT-TCC
One of the issues associated with QC-CBT-TCC is the lack of assurance that the energy obtained via QC-CBT-TCC will outperform classical algorithms in accuracy. This arises because of the approximate nature of both CBT and TCC methodologies. The TCC approach employs the standard CCSD ansatz to approximate the active space wavefunction, which implies that it does not incorporate information about higher-order excitations, such as triple and quadruple excitations, obtained from a quantum computer. Furthermore, CBT is a statistical method, and its inherent statistical nature introduces additional errors. To address this problem, we introduce a correction method as described below. The corrected energy $E_{\rm QC-CBT-TCC(c)}$ is defined as:
@@
is the CCSD energy after dressing the CCSD amplitudes but before the remaining CCSD optimization. In that case, the remaining operators $e^{\hat{T}^{\text{rest}}}$ are equivalent to the identity. Since the same error arise in $E_{\textrm{QC-CBT-TCC}}$ and $E_{\textrm{QC-CBT-TCC}}^{\rm\text{active}}$, taking the difference between these terms removes the errors in the active-space electron correlation inherent to the QC-CBT-TCC method.
Note that Izsák et al. have already employed a similar extrapolative correction method in Ref. [77] in the context of using quantum computers for quantum chemistry as well as Daniel Kats and his co-workers for FCIQMC-TCC [52].
@@ III Results and Discussion
The computational details are as follows. A chemqulacs [78] library was used to simulate quantum circuits on classical hardware. We employed disentangled unitary coupled cluster singles and doubles (UCCSD) as an ansatz for VQE and BFGS as an optimizer. CCSD, CCSD(T), CASCI, and FCI were calculated using PySCF [79]. The MR-CISD+Q method was employed to obtain reference data for the PECs of H2O and N2 because the FCI calculation was too computationally demanding for the hardware at hand. This method was implemented in the ORCA quantum chemistry program Version 5.0.3 [80]. All calculations for the PECs in this study were performed using the cc-pVDZ basis set. In Section III.3, we used the 6-31G* basis set to compute the activation energy for the Cope rearrangement. We used the canonical orbitals for all molecules except for LiH, which was calculated using the natural orbitals provided by a CCSD calculation. This, however, is an approximation because the CCSD implementation of PySCF used in this study assumes canonical orbitals, so off-diagonal terms in the Fock matrix are ignored. The active spaces were determined using the lowest energy level unoccupied orbitals and the highest energy occupied orbitals from the Hartree-Fock calculation. To transfer the CI coefficients determined by CBT to the amplitudes for CCSD, only the real parts were considered. This was necessary because usually the CCSD for a non-relativistic or non-periodic Hamiltonian expects real coefficients. The CI coefficients were normalized to ensure the normalization. Unless otherwise indicated, we applied the parameters $R=100$, $N_{\text{sample}}=10^{6}$, $N_{U}=10^{6}$, and $N_{V}=10^{6}$ for all the CBT calculations.
@@ III.1 Potential energy curves
First, we investigated the PEC of the LiH molecule. In addition to the QC-CBT-TCC(c), we show the results for QC-CBT-TCC, active space UCCSD, and FCI, as well as HF, CCSD, and CCSD(T). The selected active spaces for these calculations were two electrons and two spatial orbitals, the highest occupied and the lowest unoccupied orbitals.
Figure 2 (a) shows the PEC of the LiH molecule. All considered methods followed the qualitative features of the FCI energy curve shape, except for HF. The active space UCCSD increased the accuracy compared to that of the HF solution. This behavior is most notable in the bond dissociation region. At the equilibrium bond length, the active space UCCSD can capture some dynamic correlation. The QC-CBT-TCC and its enhanced version, the QC-CBT-TCC(c), further increased the computed energy’s accuracy. This indicates that additional dynamical correlations outside the active space are required to reach good ground-state energy. CCSD and CCSD(T) both gave energies very similar to FCI. In the high dissociation regime, where static correlation dominates, all methods except for HF gave similar results.
We focus on the energy errors of the coupled cluster and present methods compared with the FCI energy in Figure 2 (b). CCSD and CCSD(T) showed better accuracy throughout the tested bond lengths than the QC-CBT-TCC and QC-CBT-TCC(c). This can be attributed to the small static correlation within the molecule. For such a system, CCSD and CCSD(T) can capture electron correlation up to high accuracy. The energy difference between the QC-CBT-TCC and QC-CBT-TCC(c) was small. However, the QC-CBT-TCC(c) appears to produce a smoother energy error. We assume that some errors in the QC-CBT-TCC did cancel out when the QC-CBT-TCC(c) was constructed. QC-CBT-TCC and QC-CBT-TCC(c) produced energies that are lower than the FCI energy. This is possible since QC-CBT-TCC tailors the classical coupled cluster equations which are not variational.
@@
In Figure 3 (b), we show the explicit error of the presented methods compared to MR-CISD+Q. CCSD(T), and arguably CCSD, also break in the highly entangled region. QC-CBT-TCC and the QC-CBT-TCC(c) can give accurate energies. However, in the high dissociation region, QC-CBT-TCC’s error increases, whereas that of QC-CBT-TCC(c) shows this effect is reduced, indicating that the QC-CBT-TCC(c) method produces more stable energies.
@@ III.2 Number of shot dependency of QC-CBT-TCC
CBT used in our method adds a certain level of uncertainty to our computed energies. Determining the CASCI coefficients using CBT is a statistical process. To investigate this statistical behavior, we examined for the molecules LiH, H2O, and N2 the impact of varying measurement repetitions in CBT on the QC-CBT-TCC(c). We considered two constellations, one for the molecules at the equilibrium bond length [45] and one for a bond length of twice the equilibrium bond distance. Note that the geometries were not optimized. To ensure statistical significance, we repeated each measurement setup 1000 times and showed its effect on the QC-CBT-TCC(c). The parameter $R$, used to truncate the trivial computational basis in CBT, is set as 100. The behavior of LiH was similar to that of H2O and N2, and the results are shown in the Appendix. We demonstrated the effect using a box blot to demonstrate numerical data graphically.
@@
Figure 5: Energy insecurity of QC-CBT-TCC(c) for H2O with a different number of measurements for CBT. $N_{\text{sample}},N_{U}$, and $N_{V}$ were all measured in the specified quantities in the figure as the X-axis values. CBT was performed 1000 times for each setting to obtain an estimate for the statistical errors. The box indicates the range from the first to the third quartile called the inter-quartile range (IQR), with the median drawn in the middle. The box contains $50\%$ of the data. Whiskers are the lines extending 1.5 times the IQR from the first and third quantiles. Data points that exceed the whiskers are considered outliers (fliers) and are represented as single dots. (a) Number of measurements dependency at an equilibrium distance of 1.808 Bohr. (b) Number of measurements dependency at 3.617 Bohr, twice the equilibrium distance.
@@
Figure 6: Energy insecurity of QC-CBT-TCC(c) for N2 with a different number of measurements for CBT. $N_{\text{sample}},N_{U}$, and $N_{V}$ were all measured in the specified quantities in the figure as the X-axis values. CBT was performed 1000 times for each setting to obtain an estimate for the statistical errors. The box indicates the range from the first to the third quartile called the inter-quartile range (IQR), with the median drawn in the middle. The box contains $50\%$ of the data. Whiskers are the lines extending 1.5 times the IQR from the first and third quantiles. Data points that exceed the whiskers are considered outliers (fliers) and are represented as single dots. (a) Number of measurements dependency at an equilibrium distance of 2.060 Bohr. (b) Number of measurements dependency at 4.119 Bohr, twice the equilibrium distance.
To estimate the number of shots needed to reach sub-mH standard deviation of the predicted energy, we performed further calculations for the N2 at 5.355 Bohr. This system is the most challenging for CBT in our study. The other molecules need considerably fewer measurements. A standard derivation of $0.61$ mH using QC-CBT-TCC(c) can be reached using 3×3$\times$3 × $10^{7}(N_{\text{sample}}=10^{7},N_{U}=10^{7},$ and $N_{V}=10^{7})$ shots. This is in reach for today’s quantum hardware and similar to the predicted shot counts reported by Scheurer et al. [81] using the matchgate classical shadows to extract CI coefficients from a quantum computer.